# Gaussian Copula Models for Nonignorable Missing Data Using Auxiliary Marginal Quantiles

### v1

@@ 4.1 Accuracy with Sparse Auxiliary Information
Theorem 2 provides posterior consistency for the copula correlation when each $\mathcal{A}_{j}$ comprises at least $m_{j}\geq 3$ ground truth quantiles. In this section, we investigate how sensitive the contraction of the posterior is to the cardinality of $\mathcal{A}_{j}$.
@@
We simulate $2p$-dimensional observations $\{(\boldsymbol{y}_{i},\boldsymbol{r}_{i})\}_{i=1}^{n}$ for $n\in\{200,1000,5000\}$ and $p\in\{5,10,20\}$ from Gaussian copulas. For each $(n,p)$, we randomly generate $\boldsymbol{C}_{0}$ from a scaled inverse-Wishart distribution. Since each entry of $\boldsymbol{C}_{0}$ is non-zero, this generates nonignorable missingness per Lemma 1. We vary the amount of missing data by setting $\Phi(\alpha_{r_{j}})\in\{0.25,0.50\}$. Thus, marginally, missingness in each study variable is approximately 25% or 50%. For $j\leq p$, we vary $F_{j}$ so that $Y_{j}\sim\mbox{Gamma}(1,1)$ when $j\in\{1,4,7,\dots\}$; $Y_{j}\sim\mbox{t}(\nu=5,ncp=2)$ when $j\in\{2,5,8,\dots\}$; and, $Y_{j}\sim\mbox{Beta}(1,2)$ when $j\in\{3,6,9,\dots\}$. Here, $\nu$ and $ncp$ are a degrees of freedom and non-centrality parameter. Each $z_{ij}$ for $j\leq p$ is transformed to $y_{ij}$ via $F_{j}^{-1}(\Phi(\tilde{z}_{ij}))$. When $z_{i(p+j)}>0$, we set $r_{ij}=1$ and make $y_{ij}$ missing.
@@
We estimate the copula using Algorithm 1, incorporating three granularities of $\mathcal{A}$. The first includes the fully specified, true $\{F_{j}\}_{j=1}^{p}$, referred to as “Full.” The second is significantly more sparse, using just the lower/upper bounds and median, i.e., each $\mathcal{A}_{j}=\{F_{j}^{-1}(0),F_{j}^{-1}(0.5),F_{j}^{-1}(1)\}$. This is referred to as “EQL-M.” The supplement includes results for variants of EQL where $\mathcal{A}$ comprises deciles and every fourth quantile.
We also employ the EHQL, with each $\mathcal{A}_{j}=\{F_{j}^{-1}(0),F_{j}^{-1}(0.5),F_{j}^{-1}(1)\}_{j=1}^{p}$ and $s_{j}\approx 15$ intermediate quantiles. We refer to this as “EHQL-M.” The intermediate quantiles are constructed in each simulation run by first specifying 20 evenly spaced bins over the range of $\boldsymbol{y}^{obs}_{j}$ and creating $\{\mathcal{I}_{j}^{q}\}$ based on the bins occupied by the observed data. This results in approximately 15 intermediate points in each simulation run.
@@
We generate several datasets for each $(n,p,\boldsymbol{\alpha})$ setting. For each dataset and $\mathcal{A}$, we simulate 1000 posterior samples of $\boldsymbol{C}_{\boldsymbol{\theta}}$. We emphasize that fitting EQL-M and EHQL-M in these datasets does not simply parrot the data generating model, as they use only a sparse set of marginal quantiles in $\mathcal{A}$. We also include comparisons to the copula fit under the RL of Hoff, (2007), which we estimate using the sbgcop package in R. Although the RL copula is a joint model for $(\boldsymbol{y},\boldsymbol{r})$, it does not leverage any information beyond the observed data.
@@
Let $\rho_{\boldsymbol{\theta},uv}$ represent the element (correlation) in row $u$ and column $v$ of $\boldsymbol{C}_{\boldsymbol{\theta}}$. Similarly, let $\rho_{0,uv}$ be the corresponding element in the $\boldsymbol{C}_{0}$ used in data generation. Figure 3 displays 95% credible intervals based on 1000 draws of $\rho_{0,uv}-\rho_{\boldsymbol{\theta},uv}$ for the $22p(2p-1)/$ 2 unique correlation coefficients for one randomly selected dataset in the $p=5$ and 50% missingness scenario. Results for other datasets and settings are qualitatively similar; see the supplement.
@@
Theorem 2 implies that, under the EQL, $\rho_{0,uv}-\rho_{\boldsymbol{\theta},uv}$ should converge to 0 as sample size increases. This is confirmed in Figure 3 for all versions of $\mathcal{A}$. The credible intervals under EQL-M are slightly wider than those under Full, which suggests a loss in efficiency with lower levels of auxiliary information. The intervals for EHQL-M are virtually indistinguishable from those for Full. Even though EHQL-M and EQL-M leverage the same $\mathcal{A}$, EHQL-M makes greater use of the information in $\boldsymbol{y}^{obs}$ by better locating the corresponding $\boldsymbol{z}^{obs}$, which improves precision. By contrast, for the two larger values of $n$, estimates of $\boldsymbol{C}_{\boldsymbol{\theta}}$ under the RL copula are significantly biased. Finally, for $n=200$, the sampling variability is sufficiently large that inferences for all four methods are not obviously different qualitatively.
@@
Figure 3: Plots of $95\%$ credible intervals for $\rho_{0,uv}-\rho_{\boldsymbol{\theta},uv}$ for all unique correlations in $\boldsymbol{C}_{0}$ using various $\mathcal{A}$ for three sample sizes. “Full” uses full marginals; “EQL-M” uses the lower/upper bounds and median; “EHQL-M” augments each $\mathcal{A}_{j}$ for EQL-M with $s_{j}\approx 15$ estimated intermediate quantiles. Results are presented for scenario with ${}50\%$ missing values for each of $p=5$ study variables. The right-most column provides results for the copula correlation estimated under the RL. For the larger sample sizes, posterior inference for $\boldsymbol{C}_{\boldsymbol{\theta}}$ under the EQL/EHQL is accurate with minor efficiency losses when incorporating fewer auxiliary quantiles. In contrast, the inferences under RL can be biased.
@@
Figure 4: EHQL posterior samples of $\tilde{F}_{j}$ (lines) compared to true $F_{j}(y^{q}_{j})$ (triangles) and empirical CDFs (crosses) computed from $\boldsymbol{y}^{obs}$. Columns index $j=1,2,3$ from left to right. Rows index the sample size. Each $\mathcal{A}^{*}$ uses the lower/upper bounds and the median (dots) and 15 intermediate quantiles. The proposed estimator accurately outlines each true CDF and corrects the biases in the ECDF caused by nonignorable missingness.
Finally, we evaluate how closely (17) under the EHQL-M approximates each true $F_{j}$. Figure 4 displays posterior samples of $\tilde{F}_{j}(y^{q}_{j})$ for $j=1,2,3$ evaluated at the specified intermediate quantile points; it also displays the empirical CDFs (ECDF) computed from $\boldsymbol{y}^{obs}$. These ECDFs exhibit varying degrees of bias caused by the nonignorable missing data. In contrast, each $\tilde{F}_{j}(y^{q}_{j})$ accurately approximates the general shape of $F_{j}$, even though $\mathcal{A}$ only includes bounds and medians. As expected, posterior uncertainty with (17) is highest for regions with relatively small sample sizes.
@@ 4.2 Simulation of Repeated Sampling Performance
In this simulation study, we evaluate repeated sampling properties of the EHQL copula with multiple imputation inferences. We use $p=5$ variables from the North Carolina lead exposure data displayed in Table 5, namely Economically Disadvantaged (EconDisadv), Mother’s Age (mAge), Mother’s Race (mRace), an index of Neighborhood Deprivation (NDI), and end-of-fourth grade (EoG) standardized math test scores (Math_Score). These variables are mixed binary, unordered categorical, count, and continuous variables that have complex univariate and multivariate features. We collect all individuals with complete data on these five variables, excluding observations with $\lvert\text{NDI}\rvert>5$ to provide stability, which we treat as a finite population comprising approximately 165,000 individuals. Using this population, we estimate 10th, 50th, and 90th quantile regressions (Koenker, , 2010) of Math_Score on main effects of the other four variables, which we treat as population quantities. The target models aim to uncover potentially heterogeneous effects of the covariates depending on the level of academic achievement, previewing the analysis in Section 5.
@@
As auxiliary information, we assume access to selected quantiles of NDI, which we take from its marginal distribution in the constructed population. Here, we let $\mathcal{A}$ include the lower/upper bounds and median. The supplement includes results using the lower/upper bounds and 75th quantile, which are qualitatively identical to those presented here. We implement the EHQL using 15 evenly spaced quantiles across the range of observed values of NDI. For the other study variables, we use the empirical deciles in the sampled data as the quantiles in $\mathcal{A}$. We add a missingness indicator for NDI to the Gaussian copula model. We exclude the remaining indicators, which effectively models that data for the corresponding variables are MCAR.
We run the Gibbs sampler in Algorithm 1 for 5,000 iterations. After a conservative burn-in of 2,500 draws, we extract interpolated CDFs for imputation using Algorithm 2, and take the completed data in every 125th iteration to create $m=20$ multiple imputations. We estimate the targeted quantile regressions and use the combining rules of Rubin, (1987) for point estimates and 95% confidence intervals for the quantile regression coefficients. We also implement a default application of multiple imputation by chained equations (MICE) based on the mice package in R (van Buuren, , 2018). This does not use $\mathcal{A}$, although we add $R_{1}$ as a predictor for MICE. We repeat the entire procedure of sampling 5000 individuals, making missing values, and obtaining multiple imputation inferences 500 times.
Figure 5: Empirical coverage rates (left) and mean squared error (MSE) of multiple imputation point estimates (right) for the EHQL copula and MICE imputations in the repeated simulation study. For the coefficients of NDI, the EHQL provides significantly lower bias and higher coverage rates. While not shown, the average interval widths are similar.
Figure 5 summarizes the multiple imputation inferences over the 500 runs. The inferences for the coefficients subject to MCAR missingness are reasonably similar for EHQL and MICE, and generally of high quality. However, we see substantial differences among EHQL and MICE coefficient estimates for NDI. These coefficients are accurately estimated with close to nominal coverage rates under EHQL, but not MICE. Evidently, the incorporation of $\mathcal{A}$ and the flexible dependence structure under the copula model allow the imputations to reflect the multivariate relationships more accurately than MICE does. The advantage of EHQL relative to MICE is also evident in the imputations of NDI, which match the population marginal for EHQL but do not for MICE. This is visualized in Figure 6.
Figure 6: Marginal distribution of NDI in 20 completed datasets produced by EHQL and MICE. The EHQL imputations capture the bi-modality and skewness of the population marginal, whereas the MICE imputations mimic the distribution of the observed values.
@@ 5 Analysis of North Carolina Lead Exposure Data
Public health research in recent years has concluded overwhelmingly that lead exposure has adverse impacts on childhood cognitive development (e.g., Bellinger et al., , 1992; Miranda et al., , 2009; Kowal et al., , 2021; Bravo et al., , 2022). Many studies investigating this topic rely on administrative health and education datasets, linked at the individual child level, to estimate associations between lead exposure and outcomes of interest. Typically, blood-lead measurements are available only for children who are tested for lead exposure, and children are more likely to be tested when there is concern about their exposure (Kamai et al., , 2022). Consequently, abundant missingness among lead measurements is commonplace, and individuals who are measured tend to have higher levels of exposure than much of the population. Accurate imputation of lead measurements is important in this setting, as selection biases could influence population-level inferences on the associations between cognitive outcomes and lead exposure (e.g., as simulated in Section 4.2).
@@
Male .50/.50 Male infant (1 = Yes, 0 = No)
Smoker .15/.09 Mother smoked (1 = Yes, 0 = No)
NotMarried .46/.22 Not married at time of birth (1 = Yes, 0 = No)
@@
the Child Nutrition Lunch Program (1 = Yes, 0 = No)
RI .23/.18 Residential isolation index, at time of EoG test
NDI .10/-1.04 Neighborhood Deprivation Index, at time of EoG test
@@
We analyze a dataset containing information on $\sim$170,000 North Carolina children born between 2003 and 2005. The dataset is constructed by linking children’s data across three databases comprising (i) detailed birth records, which include maternal demographics, maternal and infant health measures, and maternal obstetrics history for all documented live births in N.C.; (ii) lead exposure surveillance records from a registry maintained by the state of N.C., which include integer-valued blood-lead levels; and (iii) test score data from the N.C. Education Research Data Center at Duke University, which include EoG reading and mathematics test scores as well as some demographic and socioeconomic information (CEHI, , 2020). Table 1 summarizes the variables we use, including their sample averages among children with lead exposure (Blood_lead) observed or missing. Notably, 35% of the lead measurements are missing. Other variables have missing data rates less than 0.02%.
@@
Figure 7 displays the ECDF of Blood_lead compared to three years of annual point estimates for the 50th, 75th, 90th, and 95th quantiles of lead exposure levels published by the CDC (Centers for Disease Control and Prevention, , 2022). Clearly, the observed distribution does not match the population-level quantile estimates. We therefore consider the missingness in Blood_lead as possibly missing not at random (MNAR). As we discuss later, the EHQL copula correlation estimates suggest this missingness in fact is MNAR.
@@
Figure 7: Empirical distribution of Blood_lead in the North Carolina data and population-level quantiles published by the CDC. Also displayed are average estimated quantiles of Blood_lead in completed datasets after EHQL imputations. We include CDC values consistent with the years most children were measured for lead exposure. After imputation, the marginal distribution of Blood_lead more closely matches the population-level estimates.
To adjust for this missingness, we seek to leverage the CDC estimates as auxiliary information for Blood_lead. However, the Blood_lead measurements in the North Carolina data are recorded as integers corresponding to the interval in which the child’s measurement is contained. For instance, Blood_lead=1 means that the child’s measurement was in the interval $(0,1]$ $\mu$g/ml. By contrast, the CDC measurements are continuous. In addition, the CDC estimates are national, which ignores regional or state-specific effects.
@@
We therefore use the CDC estimates to approximately locate the marginal distribution of Blood_lead and employ Algorithm 2 to infer intermediate quantiles. As evident in Figure 7, the 75th quantile estimates from the three years of CDC publications are reliably around 2.0. Because of this stability, we set $\mathcal{A}_{\text{Blood_lead}}=\{F_{\text{Blood_lead}}^{-1}(0)=0,F_{\text{Blood_lead}}^{-1}(0.75)=2,F_{\text{Blood_lead}}^{-1}(1)=10\}$. We also examine sensitivity of results when using $F_{\text{Blood_lead}}^{-1}(0.70)=2$ and $F_{\text{Blood_lead}}^{-1}(0.80)=2$. For the remaining numerical study variables, we use empirical deciles for the auxiliary quantiles. This is reasonable given the scarce missingness and large sample size. These auxiliary quantiles are not varied in the sensitivity analysis.
@@
We use the EHQL copula with the factor model in (18) to implement multiple imputation of all missing values. We include a missingness indicator for Blood_lead but not the other variables, which are almost completely observed. Their indicators comprise almost all zeros and thus are not likely to to inform the imputation but would slow computation. Consequently, the remaining study variables are treated as MCAR. For each set of auxiliary information, we estimate the EHQL copula by running Algorithm 1 for 10,000 iterations and discarding 5,000 draws as burn-in. We use every 250th posterior sample of model parameters to create $m=20$ multiple imputations. Posterior predictive checks suggest that the model reasonably describes the observed data; these are available in the supplement.
@@
Using the completed datasets, we estimate the 10th, 50th, and 90th quantile regressions of math and reading scores on main effects of all the study variables, and derive point estimates and uncertainty quantification using multiple imputation combining rules. These quantities describe potentially heterogeneous impacts of the covariates on low, middle, and high-achieving students, which is of interest to public health research (Miranda et al., , 2009).
@@
Figure 8: Multiple imputation inferences for coefficients of 10th quantile regression when Blood_lead = 2 is the 75th quantile of the marginal distribution. Compared to complete case analysis, we see sizeable differences especially for the coefficient of Blood_lead. It is estimated to be more adversely associated with EoG math scores. MICE imputations do not capture this difference for Blood_lead, as MICE mimics patterns in the observed data.
@@
Figure 8 summarizes the multiple imputation inferences for the 10th quantile regression coefficients for Blood_lead, EconDisadv, NDI, and RI using Math_Score as the response. The results are presented for $F_{\text{Blood_lead}}^{-1}(0.75)=2$; the findings are insensitive to the other values of this auxiliary quantile. We compare these inferences to those obtained by fitting the quantile regression models to complete case (CC) observations, which exclude any observations with missing study variables, and to results from $m=20$ multiple imputations using a bespoke application of MICE. Results for the other auxiliary quantile specifications, quantile regressions, and using reading scores as the response are in the supplement.
@@
After multiple imputation, we see sizeable shifts in the inferences. The estimated associations of Blood_lead, EconDisadv, and NDI with Math_Score are more adverse than in the CC analysis. The association of Blood_lead with Math_Score is estimated to be significantly stronger when using the EHQL imputations rather than the MICE imputations. Specifically, the point estimate for this coefficient under the EHQL is nearly two standard errors more negative than point estimates under CC and MICE, which is a substantively large change. The inferences under EHQL and MICE are similar for the other predictors’ coefficients, which is not unexpected as these variables have few missing values. Similar shifts are evident in the 50th and 90th quantile regressions as well.

### v2

@@ 4.1 Accuracy with Sparse Auxiliary Information
We first investigate how sensitive the contraction of the posterior is to the cardinality of $\mathcal{A}_{j}$. We simulate $2p$-dimensional observations $\{(\boldsymbol{y}_{i},\boldsymbol{r}_{i})\}_{i=1}^{n}$ for $n\in\{200,1000,5000\}$ and $p\in\{5,10,20\}$ from Gaussian copulas. For each $(n,p)$, we randomly generate $\boldsymbol{C}_{0}$ from a scaled inverse-Wishart distribution. Since each entry of $\boldsymbol{C}_{0}$ is non-zero, this generates nonignorable missingness per Lemma 1. We vary the amount of missing data by setting $\Phi(\alpha_{r_{j}})\in\{0.25,0.50\}$. Thus, marginally, missingness in each study variable is approximately 25% or 50%. For $j\leq p$, we vary $F_{j}$ so that $Y_{j}\sim\mbox{Gamma}(1,1)$ when $j\in\{1,4,7,\dots\}$; $Y_{j}\sim\mbox{t}(\nu=5,ncp=2)$ when $j\in\{2,5,8,\dots\}$; and, $Y_{j}\sim\mbox{Beta}(1,2)$ when $j\in\{3,6,9,\dots\}$. Here, $\nu$ and $ncp$ are a degrees of freedom and non-centrality parameter. Each $z_{ij}$ for $j\leq p$ is transformed to $y_{ij}$ via $F_{j}^{-1}(\Phi(\tilde{z}_{ij}))$. When $z_{i(p+j)}>0$, we set $r_{ij}=1$ and make $y_{ij}$ missing.
@@
We estimate the copula using Algorithm 1 for different granularities of $\mathcal{A}$. The first uses the fully specified, true $\{F_{j}\}_{j=1}^{p}$, referred to as “Full.” The second is significantly more sparse, using just the lower/upper bounds and median, i.e., each $\mathcal{A}_{j}=\{F_{j}^{-1}(0),F_{j}^{-1}(0.5),F_{j}^{-1}(1)\}$. This is referred to as “EQL-M.” The supplement includes results where $\mathcal{A}$ comprises deciles and every fourth quantile. We also employ the EHQL, with each $\mathcal{A}_{j}=\{F_{j}^{-1}(0),F_{j}^{-1}(0.5),F_{j}^{-1}(1)\}_{j=1}^{p}$ and $s_{j}\approx 15$ intermediate quantiles. We refer to this as “EHQL-M.” We construct intermediate quantiles in each simulation run by first specifying 20 evenly spaced bins over the range of $\boldsymbol{y}^{obs}_{j}$ and creating $\{\mathcal{I}_{j}^{q}\}$ based on the bins occupied by the observed data. This results in approximately 15 intermediate points in each simulation run.
@@
We generate several datasets for each $(n,p,\boldsymbol{\alpha})$ setting. For each dataset and $\mathcal{A}$, we simulate 1000 posterior samples of $\boldsymbol{C}_{\boldsymbol{\theta}}$. We emphasize that fitting EQL-M and EHQL-M in these datasets does not simply parrot the data generating model, as they use only a sparse set of marginal quantiles in $\mathcal{A}$. We also include comparisons to the copula fit under the RL of Hoff (2007), which we estimate using the sbgcop package in R. Although the RL copula is a joint model for $(\boldsymbol{y},\boldsymbol{r})$, it does not leverage any information beyond the observed data.
@@
Let $\rho_{\boldsymbol{\theta},uv}$ represent the element (correlation) in row $u$ and column $v$ of $\boldsymbol{C}_{\boldsymbol{\theta}}$. Similarly, let $\rho_{0,uv}$ be the corresponding element in the $\boldsymbol{C}_{0}$ used in data generation. Figure 3 displays 95% credible intervals based on 1000 draws of $\rho_{0,uv}-\rho_{\boldsymbol{\theta},uv}$ for the $22p(2p-1)/$ 2 unique correlation coefficients for one randomly selected dataset in the $p=5$ and 50% missingness scenario. Results for other datasets and settings are qualitatively similar; see Section F.1 of the supplement.
@@
Theorem 2 implies that, under the EQL, $\rho_{0,uv}-\rho_{\boldsymbol{\theta},uv}$ should converge to 0 as sample size increases regardless of the level of granularity in $\mathcal{A}$. This is confirmed in Figure 3. The credible intervals under EQL-M are slightly wider than those under Full, which suggests a loss in efficiency with lower levels of auxiliary information. However, augmenting sparse auxiliary information with intermediate quantiles (e.g., Figure 2) corrects this discrepancy; the intervals for EHQL-M are virtually indistinguishable from those for Full. Even though EHQL-M and EQL-M leverage the same $\mathcal{A}$, EHQL-M makes greater use of the information in $\boldsymbol{y}^{obs}$ by better locating the corresponding $\boldsymbol{z}^{obs}$, which improves precision. By contrast, for the two larger values of $n$, estimates of $\boldsymbol{C}_{\boldsymbol{\theta}}$ under the RL copula are significantly biased. Finally, for $n=200$, the sampling variability is sufficiently large that inferences for all four methods are not obviously different qualitatively.
@@
Figure 3: Plots of $95\%$ credible intervals for $\rho_{0,uv}-\rho_{\boldsymbol{\theta},uv}$ for all unique correlations in $\boldsymbol{C}_{0}$ using various $\mathcal{A}$ in scenario with ${}50\%$ missing values for each of $p=5$ study variables. “Full” uses full marginals; “EQL-M” uses the lower/upper bounds and median; “EHQL-M” augments each $\mathcal{A}_{j}$ for EQL-M with $s_{j}\approx 15$ intermediate quantiles. The right-most column provides results estimated under the RL. For the larger sample sizes, posterior inferences for $\boldsymbol{C}_{\boldsymbol{\theta}}$ under the EQL/EHQL are accurate with minor efficiency losses when incorporating fewer auxiliary quantiles. In contrast, the inferences under RL can be biased.
Figure 4: EHQL posterior samples of $\tilde{F}_{j}$ (lines) compared to true $F_{j}(y^{q}_{j})$ (triangles) and ECDFs (crosses) computed from $\boldsymbol{y}^{obs}$. Columns index $j=1,2,3$ from left to right. Rows index sample sizes. Each $\mathcal{A}^{*}$ uses the lower/upper bounds, the median (dots), and 15 intermediate quantiles. The EHQL accurately estimates each true CDF and corrects the biases caused by nonignorable missingness.
We next evaluate how closely (16) under the EHQL-M approximates each $F_{j}$. Figure 4 displays posterior samples of $\tilde{F}_{j}(y^{q}_{j})$ for $j=1,2,3$ evaluated at the specified intermediate quantile points; it also displays the empirical CDFs (ECDF) computed from $\boldsymbol{y}^{obs}$. These ECDFs exhibit varying degrees of bias caused by the nonignorable missing data. In contrast, each $\tilde{F}_{j}(y^{q}_{j})$ accurately approximates the general shape of $F_{j}$, even though $\mathcal{A}$ only includes bounds and medians. As expected, posterior uncertainty with (16) is highest for regions with relatively small sample sizes.
@@ 4.2 Simulation of Repeated Sampling Performance
In this section, we evaluate repeated sampling properties of the EHQL copula with multiple imputation inferences. We use $p=5$ variables from the North Carolina lead exposure data described in Section 5, namely Economically Disadvantaged (EconDisadv), Mother’s Age (MAge), Mother’s Race (MRace), an index of Neighborhood Deprivation (NDI), and end-of-fourth grade (EoG) standardized math test scores (Math_Score). These variables are mixed binary, unordered categorical, count, and continuous variables that have complex univariate and multivariate features. We collect all individuals with complete data on these five variables, excluding observations with $\lvert\text{NDI}\rvert>5$ to provide stability, which we treat as a finite population comprising approximately 165,000 individuals. Using this population, we estimate 10th, 50th, and 90th quantile regressions (Koenker, 2010) of Math_Score on main effects of the other four variables, which we treat as population quantities. The target models aim to uncover potentially heterogeneous effects of the covariates depending on the level of academic achievement, previewing the analysis in Section 5.
@@
As auxiliary information, we assume access to selected quantiles of NDI, which we take from its marginal distribution in the constructed population. Here, we let $\mathcal{A}$ include the lower/upper bounds and median. Section F of the supplement includes results using the lower/upper bounds and 75th quantile, which are qualitatively identical to those presented here, as well as for several scenarios where the median of NDI is inaccurately specified. We implement the EHQL using 15 evenly spaced quantiles across the range of observed values of NDI. For the other study variables, we use the empirical deciles in the sampled data as the quantiles in $\mathcal{A}$. We add a missingness indicator for NDI to the Gaussian copula model. We exclude the remaining indicators, which effectively models that data for the corresponding variables are MCAR.
We run the Gibbs sampler in Algorithm 1 for 5,000 iterations. After a conservative burn-in of 2,500 draws, we extract interpolated CDFs for imputation using Algorithm 2 and take the completed data in every 125th iteration to create $m=20$ multiple imputations. We estimate the targeted quantile regressions and use the combining rules of Rubin (1987) for point estimates and 95% confidence intervals for the quantile regression coefficients. We also implement a default application of multiple imputation by chained equations (MICE) based on the mice package in R (van Buuren, 2018). This does not use $\mathcal{A}$, although we add $R_{1}$ as a predictor for MICE. We repeat the entire procedure of sampling 5000 individuals, making missing values, and obtaining multiple imputation inferences 500 times.
Figure 5: Empirical coverage rates (left) and mean squared error of multiple imputation point estimates (right) for EHQL and MICE imputations in the repeated simulation study. For the coefficients of NDI, the EHQL provides significantly lower bias and higher coverage rates. While not shown, the average interval widths are similar.
Figure 5 summarizes the multiple imputation inferences over the 500 runs. The inferences for the coefficients subject to MCAR missingness are reasonably similar for EHQL and MICE, and generally of high quality. However, we see substantial differences among EHQL and MICE coefficient estimates for NDI. These coefficients are accurately estimated with close to nominal coverage rates under EHQL, but not MICE. Evidently, the incorporation of $\mathcal{A}$ and the flexible dependence structure under the copula model allow the imputations to reflect the multivariate relationships more accurately than MICE does. The advantage of EHQL relative to MICE is also evident in the imputations of NDI displayed in Figure 6, which match the population marginal for EHQL but do not for MICE.
Figure 6: Marginal distribution of NDI in 20 completed datasets produced by EHQL and MICE. EHQL imputations capture the bi-modality and skewness of the population marginal, whereas MICE imputations mimic the distribution of the observed values.
@@ 5 Analysis of North Carolina Lead Exposure Data
Public health research in recent years has concluded overwhelmingly that lead exposure has adverse impacts on childhood cognitive development (e.g., Bellinger et al., 1992; Miranda et al., 2009; Kowal et al., 2021; Bravo et al., 2022). Many studies investigating this topic rely on administrative health and education datasets, linked at the individual child level, to estimate associations between lead exposure and outcomes of interest. Typically, blood-lead measurements are available only for children who are tested for lead exposure, and children are more likely to be tested when there is concern about their exposure (Kamai et al., 2022). Consequently, abundant missingness among lead measurements is commonplace, and individuals who are measured tend to have higher levels of exposure than much of the population. Accurate imputation of lead measurements is important in this setting, as selection biases could influence population-level inferences on the associations between educational outcomes and lead exposure (e.g., as simulated in Section 4.2).
@@
Male .50/.50 Male infant (1/0 = yes/no)
Smoker .15/.09 Mother smoked (1/0 = yes/no)
NotMarried .46/.22 Not married at time of birth (1/0 = yes/no)
@@
Child Nutrition Lunch Program (1/0 = yes/no)
RI .23/.18 Residential isolation index
NDI .10/-1.04 Neighborhood deprivation index
@@
We analyze a dataset comprising $\sim$170,000 North Carolina children born between 2003 and 2005. The dataset is constructed by linking children’s data across three databases comprising (i) detailed birth records, which include maternal demographics, maternal and infant health measures, and maternal obstetrics history for all documented live births in North Carolina; (ii) lead exposure surveillance records from a registry maintained by the state, which include integer-valued blood-lead levels; and (iii) test score data from the N.C. Education Research Data Center at Duke University, which include EoG reading and mathematics test scores as well as some demographic and socioeconomic information (CEHI, 2020). Table 1 summarizes the variables we use, including their sample averages among children with lead exposure (Blood_lead) observed or missing. Notably, 35% of the lead measurements are missing. Other variables have missing data rates less than 0.02%.
@@
Figure 7 displays the ECDF of Blood_lead compared to three years of annual point estimates for the 50th, 75th, 90th, and 95th quantiles of lead exposure levels published by the CDC (Centers for Disease Control and Prevention, 2022). Clearly, the observed distribution does not match the population-level quantile estimates. We therefore consider the missingness in Blood_lead as possibly missing not at random (MNAR). As we discuss later, the EHQL copula correlation estimates suggest this missingness in fact is MNAR.
@@
Figure 7: Empirical quantiles of Blood_lead in the North Carolina data and population-level quantiles published by CDC. Also displayed are average estimated quantiles of Blood_lead in completed datasets after EHQL imputations. After imputation, the distribution of Blood_lead more closely matches the population-level distribution.
@@
To adjust for this missingness, we seek to leverage the CDC estimates as auxiliary information for Blood_lead. As evident in Figure 7, the 75th quantile estimates from three relevant years of CDC publications are reliably around 2.0. Consequently, we set $\mathcal{A}_{\text{Blood_lead}}=\{F_{\text{Blood_lead}}^{-1}(0)=0,F_{\text{Blood_lead}}^{-1}(0.75)=2,F_{\text{Blood_lead}}^{-1}(1)=10\}$. Because we do not know the distribution of lead exposure in North Carolina exactly—the CDC provides national quantiles—we examine results when using $F_{\text{Blood_lead}}^{-1}(0.70)=2$ and $F_{\text{Blood_lead}}^{-1}(0.80)=2$. This sensitivity analysis is available in Section G of the supplement. In each setting, we estimate distribution of lead exposure by setting $\{y_{\text{Blood_lead}}^{q}\}=\{3,4,5,6,7,8,9\}$ and applying Algorithm 2. For the remaining numerical study variables, we use empirical deciles for the auxiliary quantiles and the ECDF for imputation. This is reasonable given the scarce missingness and large sample size, and provides significant computational benefit for estimation of the copula model. These quantities are not varied in the sensitivity analysis.
@@
We use the EHQL copula with the factor model in (17) to implement multiple imputation of all missing values. We include a missingness indicator for Blood_lead but not the other variables, which are almost completely observed. Their indicators comprise almost all zeros and thus are not likely to to inform the imputation but would slow computation. Consequently, the remaining study variables are treated as MCAR. For each set of auxiliary information, we estimate the EHQL copula by running Algorithm 1 for 10,000 iterations and discarding 5,000 draws as burn-in. We use every 250th posterior sample of model parameters to create $m=20$ multiple imputations. Posterior predictive checks suggest that the model reasonably describes the observed data; these are available in Section G of the supplement.
@@
Using the completed datasets, we estimate the 10th, 50th, and 90th quantile regressions of math and reading scores on main effects of all the study variables, and derive point estimates and uncertainty quantification using multiple imputation combining rules. These quantities describe potentially heterogeneous impacts of the covariates on low, middle, and high-achieving students, which is of interest to public health research (Miranda et al., 2009).
@@
Figure 8: Multiple imputation inferences for coefficients of 10th quantile regression when the 75th quantile of the marginal distribution of Blood_lead = 2. Compared to complete case analysis, Blood_lead is estimated to be more adversely associated with EoG math scores. MICE imputations do not capture this difference, as MICE mimics patterns in the observed data.
@@
Figure 8 summarizes the multiple imputation inferences for the 10th quantile regression coefficients for Blood_lead, EconDisadv, NDI, and RI using Math_Score as the response. The results are presented for $F_{\text{Blood_lead}}^{-1}(0.75)=2$, and are insensitive to the other specifications of this auxiliary quantile. We compare these inferences to those obtained by fitting the quantile regression models to complete case (CC) observations, which exclude any observations with missing study variables, and to results from $m=20$ multiple imputations using a bespoke application of MICE. Results for the other auxiliary quantile specifications, quantile regressions, and using reading scores as the response are in the supplement.
@@
After multiple imputation, the associations of Blood_lead, EconDisadv, and NDI with Math_Score are more adverse than in the CC analysis. The association of Blood_lead with Math_Score is estimated to be significantly stronger when using the EHQL imputations rather than the MICE imputations. Specifically, the point estimate for this coefficient under the EHQL is nearly two standard errors more negative than point estimates under CC and MICE, which is a substantively large change. The inferences under EHQL and MICE are similar for the other predictors’ coefficients, which is not unexpected as these variables have few missing values. Similar shifts are evident in the 50th and 90th quantile regressions as well.